# Strip R-CNN: Large Strip Convolution for Remote Sensing Object Detection

### v3

@@ Abstract
While witnessed with rapid development, remote sensing object detection remains challenging for detecting high aspect ratio objects. This paper shows that large strip convolutions are good feature representation learners for remote sensing object detection and can detect objects of various aspect ratios well. Based on large strip convolutions, we build a new network architecture called Strip R-CNN, which is simple, efficient, and powerful. Unlike recent remote sensing object detectors that leverage large-kernel convolutions with square shapes, our Strip R-CNN takes advantage of sequential orthogonal large strip convolutions to capture spatial information. In addition, we enhance the localization capability of remote-sensing object detectors by decoupling the detection heads and equipping the localization head with strip convolutions to better localize the target objects. Extensive experiments on several benchmarks, for example DOTA, FAIR1M, HRSC2016, and DIOR, show that our Strip R-CNN can greatly improve previous work. In particular, our 30M model achieves 82.75% mAP on DOTA-v1.0, setting a new state-of-the-art record.
@@ 1 Introduction
Remote sensing object detection has gained significant attention in recent years due to its application in aerial images captured by drones and satellites [80, 46, 59, 33, 32, 36, 43, 37]. A popular detection pipeline is built based on rotated bounding boxes to cover objects of interest. Due to boundary discontinuity and square-like problems [73, 74, 69] and the urgent need to capture long-range information [34, 1], many research breakthroughs have been made to develop stronger rotated object detectors, including object representations [66, 68, 31, 65], IoU-simulated loss functions [73, 74, 69, 75], and foundation models [34, 53, 1], etc.
@@
To date, there are few works considering how to deal with challenging high aspect ratio objects. A generic approach widely used in previous methods is to enlarge the receptive field of models with large-kernel convolutions [12, 13, 3, 39]. A typical example should be LSKNet [34], which introduces large-kernel convolutions with a spatial selection mechanism to capture long-range contextual information. PKINet [1] further extends LSKNet and presents a multi-branch structure to enhance the detection ability for large variation of object scales. However, the parallel paradigm of leveraging multiple large-kernel convolutions exacerbates the computational burden, and the complicated block design further restricts the model efficiency. Regarding the high variation of the object aspect ratio, how to efficiently make use of large-kernel convolutions is still an open question.
@@
In this paper, we propose Strip R-CNN, which can efficiently combine the advantage of the square convolution and the strip convolution with less feature redundancy. Our design principles are two-fold. First, the new network architecture should be simple and efficient. Second, it should be good at handling objects of different aspect ratios even when they are high. Given the characteristics of the objects in remote sensing images, we propose using orthogonal large strip convolutions as the main spatial filters, which comprise the core component of our Strip R-CNN, called strip module. As such, our network is quite simple but can generalize well to even objects of especially high aspect ratios as shown in Fig. 1. Furthermore, to conquer the second challenge mentioned above, we decouple the detection heads used in previous remote sensing object detections and strengthen the localization branch with the proposed strip module. We found that this can not only assist in improving the localization capability of our method but also help more accurately regress the box angles.
@@ 2 Related Work
Remote Sensing Object Detection. Generic object detection typically relies on horizontal bounding boxes to detect objects [5, 82, 55]. However, in remote sensing scenarios, where objects are arbitrarily oriented, horizontal boxes often fail to precisely localize objects and tend to include background information or other objects [64]. Therefore, rotated bounding boxes are generally adopted for object representation. Early representations of rotated bounding boxes use five parameters $(x,y,w,h,\theta)$ [11, 71, 41]. However, due to the periodicity of the angle, training models based on this representation often face boundary discontinuity problems in regression [69, 75, 73]. To address this issue, several approaches propose improved representations [66, 65, 31, 68, 61, 76, 16, 72] for rotated bounding boxes. For instance, Oriented R-CNN [66] replaces the angle with midpoint offsets, leading to a six-parameter representation $(x,y,w,h,\triangle\alpha,\triangle\beta)$, which significantly enhances detector performance. COBB [65] introduces a continuous representation of rotated boxes with nine parameters based on the aspect ratio of the minimum enclosing rectangle to the rotated bounding box areas. There are also approaches focusing on mitigating the discontinuity problem in boundary regression through loss functions [73, 74, 26, 8, 54]. For example, GWD [73] and KLD [74] convert rotated bounding boxes into 2D Gaussian distributions and use Gaussian Wasserstein Distance and Kullback-Leibler Divergence as the loss functions. KFIoU [75] employs Gaussian modeling and Kalman filtering to propose the approximate SKewIoU Loss.
@@ 3.1 Overall Architecture
Figure 4: Structure of our basic block.
@@ 3.2 Strip Module
As discussed in Sec. 1, large square kernel convolutions provide essential long-range contextual information for remote sensing applications, but may include irrelevant features from the background regions. PKINet involves strip convolutions in model design [1] while it still relies on parallel large square kernel convolutions, which introduces computational burdens and feature redundancy. Our objective is to efficiently extract essential features for objects of varying aspect ratios. The outcome is a sequential paradigm that efficiently combines the advantages of both standard and strip convolutions without requiring additional information fusion module. In what follows, we provide a detailed description of the core of our basic block: strip module.
@@
It is important to emphasize that Strip R-CNN is much simpler than previous remote sensing object detectors using large-kernel convolutions as shown in Fig. 3. We do not utilize any spatial or channel attention mechanisms in our basic block design nor compound fusion operations with different types of large-kernel convolutions. This makes our Strip R-CNN quite simple but has great performance on different remote sensing detection benchmarks. Moreover, we found that there is no significant difference in applying horizontal strip convolutions before vertical ones or vice versa. Both approaches are effective.
@@
Figure 7: Spatial sensitivity heatmap comparison for classification and angle prediction in the output feature maps.
@@ 3.3 Detection Head with Strip Convolutions
In localization tasks, models should be sensitive to transformations, as the accuracy of localization depends on the positions of input objects. Previous strong remote sensing object detectors [1, 34] adopt the popular Oriented R-CNN framework [66], whose detection head shares the same fully connected layers for classification and localization tasks, as shown in Fig. 5. However, fully connected layers have limited spatial correlation as demonstrated in [63], making them transformation-insensitive and unsuitable for precise localization.
@@
Furthermore, predicting the parameters $x$, $y$, $w$, $h$, and $\theta$ together may lead to the issue of coupled features [78]. To address this, we decouple the prediction of $\theta$ from the other parameters ($x$, $y$, $w$, $h$). Building on this decoupling approach, we investigate the spatial sensitivity regions of the classification and angle predictions in the output feature maps. The spatial sensitivity heat map is shown in Fig. 7, where the first column represents the classification sensitivity, and the second shows the angle sensitivity. Our observation reveals that the spatial sensitivity for classification is concentrated in the central areas of the objects, while the sensitivity for angle is primarily focused near the object borders. The sensitive regions for classification and angle predictions share some overlap and exhibit some complementary characteristics. Based on this, we propose the use of shared fully connected layers for both classification and angle predictions. The structure of the strip head is shown at the bottom of  Fig. 5. Detailed configurations are described below.
@@
Angle head: For angle prediction, we adopt three fully connected layers to effectively estimate angular information, which we found works well. Note that the first two fully connected layers share parameters with the classification head.
Loss function: All three heads, including classification head, localization head, and angle estimation head, are jointly trained end-to-end. The overall loss function can be written as follows:
$\mathcal{L}=\mathcal{L}_{c}+\mathcal{L}_{l}+\mathcal{L}_{a},$ (2)
where $\mathcal{L}_{c}$ is cross entropy loss for classification head, $\mathcal{L}_{l}$, and $\mathcal{L}_{a}$ are the Smooth L1 loss [17] for localization head and angle head, respectively.
It is worth noting that our detection head incorporates tce is cross entropy loss, proposed strip module. This design largely improves the detection capability compared to the original Oriented R-CNN head. As shown in Fig. 6, from the visualization of the spatial sensitivity of different methods, we can see that the Oriented R-CNN head has a similar spatial correlation pattern on the output feature maps. However, our strip head has much more spatial correlations in the output feature maps. We will show that this enables our method to be able to better localize objects even of higher aspect ratios.
@@ 4 Experiments
4.1 Datasets
We conduct extensive experiments on five popular remote sensing object detection datasets.
@@ 4.1 Datasets
  • FAIR1M-v1.0 [58] is a remote sensing dataset that consists of 15,266 images, 1 million instances, and 37 categories.
@@
4.2 Implementation Details
Our training process is divided into two stages: pretraining on ImageNet [10] followed by fine-tuning on downstream remote sensing datasets. In the ablation experiments, we consistently train all the models for 100 epochs. To achieve higher performance, we train the models for 300 epochs in the final results table. The number of training epochs for the DOTA, DOTAv1.5, HRSC2016, FAIR1M-v1.0, and DIOR-R datasets are set to 12, 12, 36, 12, and 12, respectively, following previous methods. The learning rates are set to 0.0001, 0.0001, 0.0004, 0.0001, and 0.0001, respectively. The input sizes for the HRSC2016 and DIOR-R datasets are $800\times 800$, while for the DOTA-v1.0, DOTA-v1.5 and FAIR1M-v1.0 datasets, the input sizes are $1024\times 1024$. During training, we employ the AdamW [44] optimizer with $\beta_{1}=0.9,\beta_{2}=0.999$, and a weight decay of 0.05. All the models are trained on 8 NVIDIA 3090 GPUs with a batch size of 8, and test is conducted on a single NVIDIA 3090 GPU.
@@ 4.3 Main Results
Table 2: Comparisons with different backbone models on the DOTA-v1.0 dataset. Params and FLOPs are computed for backbone only. All backbones are pretrained on ImageNet for 300 epochs. Our StripNet-S achieves higher mAP than previous popular backbones.
@@
Results on DOTA-v1.0 [64]. We conduct a comparative analysis of different models and present detailed results for mean Average Precision (mAP) and Average Precision (AP) across categories on the DOTA dataset. As shown in Tab. 3, our single-scale evaluation demonstrates a 1.67% improvement over PKINet-S. Furthermore, with multi-scale training and testing, we achieve 82.28% mAP for a single model. By ensembling the results of RTMDet and Strip R-CNN, following the model ensemble strategy in MoCAE [51], we further achieve 82.75% mAP, setting a new state-of-the-art record.
@@
Table 7: Comparisons with the SOTA methods on the DIOR-R Datasets. The StripNet-S backbone is pretrained on ImageNet for 300 epochs.
@@ 4.4 Ablation Studies
Kernel size of strip convolutions. The kernel size in strip convolutions is critical for our proposed strip module. We experiment with large kernel sizes, starting from 11 and increasing in increments of 4. Smaller kernels are ineffective in capturing features of high aspect ratio objects, while a kernel size of 15 yielded satisfactory results, with further improvement observed at 19. We also test various kernel sizes at different network stages, exploring both increasing and decreasing strategies. Larger kernels in shallow layers combined with smaller kernels in deeper layers produce good results. In contrast, using smaller kernels in the shallow layers leads to significant performance drops, indicating that shallow layers benefit from larger receptive fields. Consequently, we select 19 as the optimal kernel size for the strip module at all stages.
@@
Ablation on strip module design We conduct ablation experiments to analyze the design choices in the strip module. First, we assess the role of depth-wise square convolution. Removing this component leads to a significant performance drop, emphasizing the importance of square convolution for capturing features of square-shaped objects. Next, we examine the integration of horizontal and vertical large strip convolutions, comparing parallel and sequential arrangements. The sequential configuration outperforms the parallel one, as the latter lacks effective two-dimensional modeling, merely combining the two strip convolutions without capturing the overall object structure. Furthermore, substituting the sequential large strip convolutions with either a $19\times 19$ large kernel convolution or a $7\times 7$ dilated convolution with a dilation rate of 3 results in noticeable performance loss, further validating the effectiveness of the large strip convolutions. The results are reported in Tab. 9.
@@
Table 10: Ablation study of the strip head design. fc refers to using two fully connected layers in this branch. conv refers to using two depth-wise $3\times 3$ convolutional layers followed by point-wise convolution in this branch. strip refers to replacing the latter $3\times 3$ convolutional layer with our designed strip module. We adopt the StripNet-S backbone pretrained on ImageNet for 100 epochs.
@@
       $(x,y,w,h,\theta)_{fc}$        81.75
       $(x,y)_{conv},(w,h)_{conv},(\theta)_{fc}$        81.76
       $(x,y)_{conv},(w,h)_{conv},(\theta)_{conv}$        81.72
       $(x,y)_{strip},(w,h)_{conv},(\theta)_{fc}$        81.88
       $(x,y)_{conv},(w,h)_{strip},(\theta)_{fc}$        81.77
       $(x,y)_{conv},(w,h)_{conv},(\theta)_{strip}$        81.81
       $(x,y,w,h)_{strip},(\theta)_{fc}$        81.96
       $(x,y,w,h)_{strip},(\theta)_{strip}$        81.93
The effectiveness of the strip head. As can be seen in Tab. 11, we compare the performance of two different detectors before and after incorporating the strip head. The results demonstrate that our strip head consistently improves the performance of other detectors, confirming its effectiveness and generalizability. For the ROI Transformer, our method achieves a significant improvement of 5.0 mAP, which is a remarkable increase. Similarly, the other detector Rotated Faster R-CNN also receives performance gains, further validating that our approach is compatible with a wide range of detectors and is able to reliably enhance detection performance.

### v4

@@ Abstract
While witnessed with rapid development, remote sensing object detection remains challenging for detecting high aspect ratio objects. This paper shows that large strip convolutions are good feature representation learners for remote sensing object detection and can detect objects of various aspect ratios well. Based on large strip convolutions, we build a new network architecture called Strip R-CNN, which is simple, efficient, and powerful. Unlike recent remote sensing object detectors that leverage large-kernel convolutions with square shapes, our Strip R-CNN takes advantage of sequential orthogonal large strip convolutions in our backbone network StripNet to capture spatial information. In addition, we improve the localization capability of remote-sensing object detectors by decoupling the detection heads and equipping the localization branch with strip convolutions in our strip head. Extensive experiments on several benchmarks, for example DOTA, FAIR1M, HRSC2016, and DIOR, show that our Strip R-CNN can greatly improve previous work. In particular, our 30M model achieves 82.75% mAP on DOTA-v1.0, setting a new state-of-the-art record. Our code will be made publicly available.
@@ 1 Introduction
Remote sensing object detection has gained significant attention in recent years due to its application in aerial images captured by drones and satellites [78, 44, 57, 31, 30, 34, 41, 35]. A popular pipeline is built on the basis of rotated boxes to cover objects of interest. Due to boundary discontinuity and square-like problems [71, 72, 67] and the urgent need to capture long-range information [32, 1], many research breakthroughs have been made to develop stronger rotated object detectors, including object representations [64, 66, 29, 63], IoU-simulated loss functions [71, 72, 67, 73], and foundation models [32, 51, 1], etc.
@@
To date, there are few works considering how to deal with challenging high aspect ratio objects. A generic approach widely used in previous methods is to enlarge the receptive field of models with large-kernel convolutions [11, 12, 3, 37]. A typical example should be LSKNet [32], which introduces large-kernel convolutions with a spatial selection mechanism to capture long-range contextual information. PKINet [1] further extends LSKNet with a parallel large square conv structure to enhance the performance for large variation of object scales. However, the parallel paradigm of using multiple large-kernel convolutions exacerbates the computational burden, and the complicated block design further restricts model efficiency. For the high variation of the object aspect ratio, how to efficiently use large-kernel convolutions is still an open question.
@@
In this paper, we propose Strip R-CNN, which can efficiently combine the advantage of the square convolution and the strip convolution with less feature redundancy. Our design principles are two-fold. First, the new network architecture should be simple and efficient. Second, it should be good at handling objects of different aspect ratios even when they are high. Given the characteristics of the objects in remote sensing images, we propose using orthogonal large strip convolutions as the main spatial filters, which comprise the core component of our StripNet backbone, called strip module. As such, our network is quite simple but can generalize well to even objects of especially high aspect ratios as shown in Fig. 1. Furthermore, to conquer the second challenge mentioned above, we decouple the detection heads used in previous remote sensing object detections and strengthen the localization branch with strip convolutions in our strip head. We found that this can not only assist in improving the localization capability of our method but also help more accurately regress the box angles.
@@ 2 Related Work
Remote Sensing Object Detection. In remote sensing scenarios, where objects are arbitrarily oriented, rotated bounding boxes are generally adopted for object representation. Early representations of rotated bounding boxes use five parameters $(x,y,w,h,\theta)$ [10, 69, 39, 77]. However, due to the periodicity of the angle, training models based on this representation often face boundary discontinuity problems in regression [67, 73, 71]. To address this issue, several approaches propose improved representations [64, 63, 29, 66, 59, 74, 15, 70] for rotated bounding boxes. For instance, Oriented R-CNN [64] replaces the angle with midpoint offsets, leading to a six-parameter representation $(x,y,w,h,\triangle\alpha,\triangle\beta)$, which significantly enhances detector performance. COBB [63] introduces a continuous representation of rotated boxes with nine parameters based on the aspect ratio of the minimum enclosing rectangle to the rotated bounding box areas. There are also approaches focusing on mitigating the discontinuity problem in boundary regression through loss functions [71, 72, 24, 7, 52]. For example, GWD [71] and KLD [72] convert rotated bounding boxes into 2D Gaussian distributions and use Gaussian Wasserstein Distance and Kullback-Leibler Divergence as the loss functions. KFIoU [73] uses Gaussian modeling and Kalman filtering and propose the SKewIoU Loss.
@@ 3.1 Overall Architecture
Figure 4: Structure of our basic block of StripNet backbone.
@@ 3.2 Strip Module
As discussed in Sec. 1, large square kernel convolutions provide essential long-range contextual information for remote sensing applications. PKINet [1] introduces parallel large square kernel convolutions and spatial attention mechanism to extract spatial information. However, the parallel paradigm of leveraging multiple large-kernel convolutions exacerbates the computational burden, and the complicated block design makes the model not efficient. Our objective is to efficiently extract essential features for objects of varying aspect ratios. The outcome is a sequential paradigm that efficiently combines the advantages of both standard and strip convolutions without requiring additional information fusion module. In what follows, we provide a detailed description of the core of our basic block: strip module.
@@
It is important to emphasize that our backbone network StripNet is much simpler than previous remote sensing object detectors using large-kernel convolutions as shown in Fig. 3. We do not utilize any spatial or channel attention mechanisms in our basic block design nor compound fusion operations with different types of large-kernel convolutions. This makes our StripNet quite simple but has great performance on different remote sensing detection benchmarks. Moreover, we found that there is no significant difference in applying horizontal strip convolutions before vertical ones or vice versa. Both approaches are effective.
@@
Figure 7: Spatial sensitivity heatmap comparison for classification, localization and angle prediction in the output feature maps.
@@ 3.3 Detection Head with Strip Convolutions
In localization tasks, models should be sensitive to transformations, as the accuracy of localization depends on the positions of input objects. Previous strong rotated object detectors [1, 32] adopt the Oriented R-CNN framework [64], whose detection head shares the same linear layers for classification and localization, as shown in Fig. 5. However, linear layers have limited spatial correlation as demonstrated in [61], making them transformation-insensitive and unsuitable for precise localization.
@@
Furthermore, predicting the parameters $x$, $y$, $w$, $h$, and $\theta$ together may lead to the issue of coupled features [76]. To address this, we decouple the prediction of $\theta$ from the other parameters ($x$, $y$, $w$, $h$). Building on this decoupling approach, we investigate the spatial sensitivity regions of the classification and angle predictions in the output feature maps. The spatial sensitivity heat map is shown in Fig. 7, where the first column represents the classification sensitivity, and the second shows the angle sensitivity. Our observation reveals that the spatial sensitivity for classification is concentrated in the central areas of the objects, while the sensitivity for angle is primarily focused near the object borders. The sensitive regions for classification and angle predictions share some overlap and exhibit some complementary characteristics. Based on this, we propose the use of shared fully connected layers for both classification and angle predictions. The structure of the strip head is shown at the bottom of  Fig. 5. Detailed settings are described below.
@@
Angle head: For angle prediction, we adopt three fully connected layers to estimate angular information, which we found works well. Note that the first two fully connected layers share parameters with the classification head.
It is worth noting that our detection head incorporates the proposed strip module. This design largely improves the detection capability compared to the original Oriented R-CNN head. As shown in Fig. 6, from the visualization of the spatial sensitivity of different methods, we can see that the Oriented R-CNN head has a similar spatial correlation pattern on the output feature maps. However, our strip head has much more spatial correlations in the output feature maps. We will show that this enables our method to be able to better localize objects even of higher aspect ratios.
@@ 4 Experiments
4.1 Experiment Setup
Datasets. We conduct extensive experiments on five popular remote sensing object detection datasets.
@@ 4.1 Experiment Setup
  • FAIR1M-v1.0 [56] is a remote sensing dataset consisting of 15,266 images, 1 million instances, and 37 categories.
@@
Implementation details. Our training process is divided into two stages: pretraining on ImageNet [9] and fine-tuning on downstream remote sensing datasets. For ablation experiments, we train all the models for 100 epochs. To achieve higher performance, we train the models for 300 epochs for the final results. The number of training epochs for DOTA, DOTAv1.5, HRSC2016, FAIR1M-v1.0, and DIOR-R are set to 12, 12, 36, 12, and 12, respectively, following previous methods. Learning rates are set to 0.0001, 0.0001, 0.0004, 0.0001, and 0.0001, respectively. The input sizes for HRSC2016 and DIOR-R are $800\times 800$, while for the DOTA-v1.0, DOTA-v1.5 and FAIR1M-v1.0 datasets, the input sizes are $1024\times 1024$. During training, we employ the AdamW [42] optimizer with $\beta_{1}=0.9,\beta_{2}=0.999$, and a weight decay of 0.05. All the models are trained on 8 NVIDIA 3090 GPUs with a batch size of 8, and test is conducted on a single NVIDIA 3090 GPU.
@@ 4.2 Main Results
Table 2: Comparisons with different backbone models on DOTA-v1.0. Params and FLOPs are computed for backbone only. All backbones are pretrained on ImageNet for 300 epochs. Our StripNet-S achieves higher mAP than previous popular backbones.
@@
Results on DOTA-v1.0 [62]. We conduct a comparative analysis of different models and present detailed results for mean Average Precision (mAP) and Average Precision (AP) across categories on the DOTA dataset(refer to more model results in Supplementary Materials). As shown in Tab. 3, our single-scale evaluation demonstrates a 1.67% improvement over PKINet-S. Furthermore, with multi-scale training and testing, we achieve 82.28% mAP for a single model. By ensembling the results of RTMDet and Strip R-CNN, following the model ensemble strategy in MoCAE [49], we achieve 82.75% mAP, setting a new state-of-the-art record.
@@
Table 7: Comparisons with SOTA methods on DIOR-R. The StripNet-S backbone is pretrained on ImageNet for 300 epochs.
@@ 4.3 Ablation Studies
Kernel size of strip convolutions. The kernel size in strip convolutions is critical for our proposed strip module. We experiment with large kernel sizes, starting from 11 and increasing in increments of 4. Smaller kernels are ineffective in capturing features of high aspect ratio objects, while a kernel size of 15 yielded satisfactory results, with further improvement observed at 19. To further show the advantage of large strip convolution, we add a visual analysis of the feature representations of different kernel sizes. Fig. 8 shows that strip convs with a larger kernel size can learn the features of slender objects with more precise localization information and well sharpen the object boundaries. We also test various kernel sizes at four stages in our StripNet backbone, exploring both increasing and decreasing strategies. Larger kernels in shallow layers combined with smaller kernels in deeper layers produce good results. In contrast, using smaller kernels in the shallow layers leads to significant performance drops, indicating that shallow layers benefit from larger receptive fields. Consequently, we select 19 as the optimal kernel size for the strip module at all stages of StripNet backbone.
Figure 8: Feature map visulization of different kernel sizes. Kernel size 19 leads to more accurate features of high aspect ratio objects.
@@
Ablation on strip module design. We perform ablation experiments to analyze the design choices in the strip module as shown in Tab. 9. First, we assess the role of depth-wise square convolution. Removing this component leads to a large performance drop, emphasizing the importance of square convolution for capturing features of square-shaped objects. Next, we examine the integration of horizontal and vertical large strip convolutions, comparing parallel and sequential arrangements. The sequential configuration outperforms the parallel one, as the latter lacks effective two-dimensional modeling, merely combining the two strip convolutions without capturing the overall object structure. Furthermore, substituting the sequential large strip convolutions with either a $19\times 19$ large kernel convolution or a $7\times 7$ dilated convolution with a dilation rate of 3 results in noticeable performance loss, further validating the effectiveness of the large strip convolutions.
[…]
RoI Trans. [10] ResNet-50 [21] 41.1M 211.4G 74.61
StripNet-S 13.3M 52.3G 81.72
S2ANet [19] ResNet-50 [21] 41.1M 211.4G 79.42
StripNet-S 13.3M 52.3G 80.52
R3Det [68] ResNet-50 [21] 41.1M 211.4G 76.47
StripNet-S 13.3M 52.3G 79.01
YOLO [43] CSPNeXt-l [21] 52.3M 80.2G 78.02
StripNet-S 13.3M 52.3G 78.50
ARS-DETR [79] ResNet-50 [21] 41.1M 211.4G 72.08
StripNet-S 13.3M 52.3G 75.07
Effectiveness of the StripNet backbone. To validate the generality and effectiveness of our proposed StripNet backbone, we conduct experiments using various remote sensing detection frameworks. As can be seen in Tab. 17, Our method improves RoI Transformer [10] by 7.11%. It also enhances S2ANet and R3Det by 0.9% and 2.54%, respectively. When integrated into YOLO and DETR frameworks, our approach consistently boosts performance, demonstrating its broad compatibility and effectiveness in enhancing various detectors.
@@
Table 11: Ablation study of the strip head design. fc refers to using two fully connected layers. conv refers to using two $3\times 3$ convolutional layers. strip refers to replacing the latter $3\times 3$ convolutional layer with our designed strip module. We adopt the StripNet-S backbone pretrained on ImageNet for 100 epochs.
@@
      $(x,y,w,h,cls,\theta)_{fc}$       81.75
      $(x,y)_{conv},(w,h)_{conv},(cls)_{fc},(\theta)_{fc}$       81.76
      $(x,y)_{conv},(w,h)_{conv},(cls)_{fc},(\theta)_{conv}$       81.72
      $(x,y)_{strip},(w,h)_{conv},(cls)_{fc},(\theta)_{fc}$       81.88
      $(x,y)_{conv},(w,h)_{strip},(cls)_{fc},(\theta)_{fc}$       81.77
      $(x,y)_{conv},(w,h)_{conv},(cls)_{fc},(\theta)_{strip}$       81.81
      $(x,y,w,h)_{strip},(cls)_{fc},(\theta)_{fc}$       81.86
      $(x,y,w,h)_{strip},(cls,\theta)_{fc}$       81.96
      $(x,y,w,h)_{strip},(cls)_{fc},(\theta)_{strip}$       81.93
Effectiveness of the strip head. As can be seen in Tab. 12, we compare the performance of two different detectors before and after incorporating the strip head. The results demonstrate that our strip head consistently improves the performance of other detectors, confirming its effectiveness and generalizability. For the ROI Transformer, our method achieves a significant improvement of 5.0 mAP, which is a remarkable increase. Similarly, the other detector Rotated Faster R-CNN also receives performance gains, further validating that our approach can be used for a wide range of detectors and is able to enhance their detection performance.
Table 12: Effectiveness of strip head on other rotated detectors.